3 metalliic cubes of dimensions 6 meter, 8meters, 10 meters are melted to form a single big cube/ find the dimensions of the cube
Answer by mananth(16946)

Volume of cube I = 6^3= 216 cu.m
Volume of cube II = 8^3 = 512 cu.m
Volume of cube III = 10^3= 1000 cu.m
Add up
Total volume = 1728 cu.m
Volume of new cube = 1728 cu.m
length of side = cube root of 1728 = 12
side = 12m